A beaker with 2.00×102 mLmL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 MM. A student adds 4.20 mLmL of a 0.410 MM HClHCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.740.
Answer #1

The calculations of this question uploaded as images.Total imolarity of acid and conjugate base in buffer Sollution is 0.1 molar la Let us suppose concentration of acid is a (M)08 dog love) = (5 - 4.790) = 0.26 0 00 Oll-x = 10 y 10 - 1.8197 = 1.8197 -1 = 18197 03.07. Ou = (1.5197 +i) = 2.8194 67, X =Volume of buffer solution = 200 me. 200 del 0.0554 (M) CalcooH = 20000 : 0354 mol Cty cooh 50.0070840 f HyCOOH, 4.2 ml of 0.42004 0.06 4 c (м) и е с 2 а х 0.0 646 back 100 оf еңгі - 0.012 2 4d of eц сов Studio».ly, asuut of c, tо - о 1202 - 6.00 (122